# /csteps/fill color=green!10

### v1

@@ I Introduction
Hiding-based countermeasures, such as shuffling, on the other hand, can greatly increase the attacker’s effort while keeping the overhead manageable even in embedded systems [2]. There have been several proposals to use shuffling for neural network implementations. Nozaki and Yoshikawa [5], and Brosch et al. [6] proposed a software-based shuffling of multiplications within a neuron to make the parameter recovery infeasible. The random shuffle algorithm that is commonly used in such cases is the Fisher-Yates algorithm [7] as it provides optimal $\mathcal{O}(n)$ time complexity and produces an unbiased permutation. However, it was shown by Ganesan et al. [8] that this shuffling method can be easily broken by an SCA targeting the division operation. They provided a way to reorder the parts of SCA traces back to the original order, thus allowing the parameter recovery. To counteract this attack, they proposed to realize the shuffling in hardware, avoiding the usage of the Fisher-Yates algorithm. While their method is secure and efficient, it requires an additional hardware implementation which is not an option for general-purpose microcontrollers. In this paper, we aim to overcome this limitation by developing a side-channel secure way to realize shuffling in software.
@@ 9                   $r_{e}[\texttt{exponent}]=r[j,t]$
To ensure the completeness of our evaluation, in Subsection IV-A, we will outline the CPA attack methodology for recovering the different components of a secret weight value, ultimately reconstructing the entire weight. Subsequently, in Subsection IV-B, we will demonstrate how the different components of a secret weight value can be successfully recovered in an unprotected implementation, as well as how our proposed countermeasure effectively mitigates such an attack.
@@ IV-A CPA Attack Steps
IV-B Results
@@
The power traces for the first hidden layer computations are illustrated in Figure 2 for both the unprotected network and the network with shuffled multiplications. As shown, the computations for the five neurons are clearly distinguishable in both cases, marked by the red dotted lines.
These observations indicate that recovering the network architecture (e.g., the number of neurons per layer) through visual inspection of the power traces remains feasible in both scenarios. This is expected as our countermeasure is not designed to conceal the network architecture but rather to protect the secret weight parameters from being extracted using CPA.
@@
[…]
The weight values are randomly generated within $[-2,2]$ with a precision of $0.01$, resulting in the set:
@@
The CPA attack results for the unprotected and protected implementations are presented in Figures 4 and 5 respectively, where the absolute correlations for the correct values are highlighted in red, while those for incorrect values are shown in gray. For the first byte of the mantissa, only the maximum absolute correlation among incorrect values for each time sample is plotted due to the large number of possible values.
For the attacks on the unprotected implementation, Figure 4 clearly demonstrates that the correct weight value can be successfully recovered.
For attacks on the protected implementation, Figure 5(a) shows that the sign bit of the weight can still be recovered. We believe the reason lies in the composition of the network’s weights and inputs. Among the seven weights used, four are positive. Additionally, the inputs to other neurons were set to $0.5$, increasing the likelihood of intermediate values being positive. Consequently, the sign bit value of 00 exhibits higher absolute correlations. We would like to note that recovering the sign bit itself does not significantly reduce the search complexity of the entire floating point value. For other components of the weight value, the peaks corresponding to the correct values are significantly lower compared to those of incorrect values.
@@ V Discussion
Figure 6: Computation overhead of the proposed protected single-layer computation compared to standard shuffled single-layer computation. The x-axis represents the number of input and output neurons, which are assumed to be equal.
@@ V-A Overheads
Compared to the original Fisher-Yates shuffling method, the masked shuffling introduces an overhead of 3.38×3.38$\times$3.38 × for $N=1000$ (i.e., with $1000$ input neurons) and 2.48×2.48$\times$2.48 × for $N=100$.
For a single layer’s shuffled computation, incorporating our proposed shuffling operation (Algorithm 3) introduces an overhead of $4\%$ compared to the standard shuffled computation using the original Fisher-Yates algorithm (Algorithm 1) when the layer has $100$ input neurons and $100$ output neurons, with ReLU as the activation function. When the number of input and output neurons increases to $1000$, the overhead decreases to $0.49\%$. A plot illustrating the overhead as a function of the number of neurons (assuming an equal number of input and output neurons) is provided in Figure 6.
The memory overhead compared to Fisher-Yates shuffling-based countermeasures arises from storing the two secret arrays, $S_{1}$ and $S_{2}$. As previously noted, the size of $S_{1}$ and $S_{2}$ is determined by the maximum number of neurons in any layer of the network. Specifically, their size is equal to this maximum number minus two.
@@ VI Conclusion
For the future work, it would be interesting to evaluate a combination of different hiding-based countermeasures, such as shuffling and desynchronization, or a combination of hiding and masking.

### v2

@@ I Introduction
Hiding-based countermeasures, such as shuffling, on the other hand, can greatly increase the attacker’s effort while keeping the overhead manageable even in embedded systems [2]. Shuffling has been extensively evaluated by the cryptography community, for example, [5] provides an analysis of its application. Herbst et al. [6] proposed it for shuffling Sboxes and key additions in a hardware AES implementation. The most recent paper by Xu et al. [7] modifies the Fisher-Yates algorithm for hardware usage to show no TVLA leakage with $10^{7}$ electromagnetic traces, measuring the post-quantum Kyber algorithm.
There have been several proposals to use shuffling for neural network implementations. Nozaki and Yoshikawa [8], and Brosch et al. [9] proposed a software-based shuffling of multiplications within a neuron to make the parameter recovery infeasible. The random shuffle algorithm that is commonly used in such cases is the Fisher-Yates algorithm [10] as it provides optimal $\mathcal{O}(n)$ time complexity and produces an unbiased permutation. However, it was shown by Ganesan et al. [11] that this shuffling method can be easily broken by an SCA targeting the division operation. They provided a way to reorder the parts of SCA traces back to the original order, thus allowing the parameter recovery. To counteract this attack, they proposed to realize the shuffling in hardware, avoiding the usage of the Fisher-Yates algorithm. While their method is secure and efficient, it requires an additional hardware implementation which is not an option for general-purpose microcontrollers. In this paper, we aim to overcome this limitation by developing a side-channel secure way to realize shuffling in software.
@@ 9                   $r_{e}[\texttt{exponent}]=r[j,t]$
To ensure the completeness of our evaluation, in Subsection IV-A, we will outline the CPA attack methodology for recovering the different components of a secret weight value, ultimately reconstructing the entire weight. Subsequently, in Subsection IV-C, we will demonstrate how the different components of a secret weight value can be successfully recovered in an unprotected implementation, as well as how our proposed countermeasure effectively mitigates such an attack.
@@ IV-A CPA Attack Steps
IV-B Experimental Setup
We used ChipWhisperer Lite with ARM Cortex-M4 (STM32F3) target as the basis for our setup. The target frequency was set to $7.3728$ MHz, and the ADC frequency was set to 4×4$\times$4 × that. The value shown in the y-axis of the power traces is proportional to the current going through the shunt resistor — we refer to this as “leakage” in the graphs as these are not exact current values.
IV-C Results
@@
The power traces for the first hidden layer computations are illustrated in Figure 2 for both the unprotected network and the network with shuffled multiplications. As shown, the computations for the five neurons are clearly distinguishable in both cases, marked by the red dotted lines. The visual inspection allowing to determine the number of layers and neurons per layer has been described in [3]. As the patterns for multiplications and activation functions are clearly distinguishable from each other, this allows the attacker to split the power trace into segments, each containing one layer. Within the layer trace segment, the attacker can then zoom in to identify the repeating multiplication patterns.
These observations indicate that recovering the network architecture (e.g., the number of neurons per layer) through visual inspection of the power traces remains feasible in both scenarios. This is expected as our countermeasure is not designed to conceal the network architecture but rather to protect the secret weight parameters from being extracted using CPA. We would like to note that this allows a stronger attacker model.
@@
The CPA attack results for the unprotected and protected implementations are presented in Figures 4 and 5 respectively, where the absolute correlations for the correct values are highlighted in red, while those for incorrect values are shown in gray. For the first byte of the mantissa, only the maximum absolute correlation among incorrect values for each time sample is plotted due to the large number of possible values [^2] Since the weight values are constrained to the range $[-2,2]$ with two decimal places of precision, not all possible values for the exponent and mantissa bytes can be represented, leading to variations in the number of incorrect hypotheses plotted..
For the attacks on the unprotected implementation, Figure 4 clearly demonstrates that the correct weight value can be successfully recovered. As detailed in Algorithm 4, the absolute correlation values were grouped differently depending on the specific weight component being targeted. For each time sample, the highest absolute correlation within each group was selected and plotted. Notably, all red lines in the figures exhibit the same behavior, indicating that the true weight value consistently achieves the highest absolute correlation, regardless of how the correlation coefficients are grouped. Furthermore, the minimal differences between the red and gray lines in Figure 4 (a) highlight the importance of recovering each weight component separately. In fact, if we plot the absolute correlation for the true weight alongside the highest absolute correlation among incorrect weight hypotheses, the resulting plot closely resembles Figure 4 (a), making it challenging to distinguish the true weight from the incorrect ones.
For attacks on the protected implementation, Figure 5(a) shows that the sign bit of the weight can still be recovered. Unlike Figure 4(a), where the exact location of the target multiplication is known for each execution, in this case, the precise timing is unknown. As a result, all seven multiplications were analyzed for each execution. Among these, only one multiplication varies per execution, while the remaining six involve three fixed positive weights, three fixed negative weights, and six fixed positive inputs (0.5). Since the majority of these values are positive and appear randomly within the analyzed interval, we believe this explains why the positive sign exhibits a higher absolute correlation with the leakage. We would like to note that recovering the sign bit itself does not significantly reduce the search complexity of the entire floating point value. For other components of the weight value, the peaks corresponding to the correct values are significantly lower compared to those of incorrect values.
@@
IV-D Impact of Increasing the Number of Traces on Attack Performance
To assess how the attack evolves as the number of traces increases, we conducted the same attack using varying numbers of traces for both protected and unprotected implementations.
Figures 6 and 7 illustrate the results for the unprotected and protected implementations, respectively. We collected a total of $5,000$ traces for the unprotected implementation and $50,000$ traces for the protected implementation. To evaluate the attack’s effectiveness, we varied the number of traces from $3$ to $5,000$ for the unprotected case and up to $50,000$ for the protected, using an appropriate step size. For each trace count, we recorded the highest absolute correlation coefficient associated with the correct weight components and the maximum absolute correlation corresponding to incorrect weight components.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 6: CPA attack results with different numbers of traces for the unprotected implementation. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $1.43$, while the gray lines correspond to incorrect values.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 7: CPA attack results with different numbers of traces for the protected implementation. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $1.43$, while the gray lines correspond to incorrect values.
We observe that at lower trace counts, both correct and incorrect weight values exhibit high absolute correlation values in both protected and unprotected implementations. This behavior can be attributed to the limited number of data points used in the Pearson correlation calculation. Since the correlation coefficient is derived from sample means and standard deviations, small datasets result in unstable estimates, leading to significant fluctuations in correlation values. As a consequence, high correlations may occur due to random variations. We note that a similar phenomenon has been reported in [9].
As the number of traces increases, the correlation stabilizes as statistical noise averages out. The results in Figure 6 demonstrate that in the unprotected implementation, the attack achieves success with only a few hundred traces, effectively distinguishing the correct weight components from incorrect ones. Conversely, in the protected implementation (Figure 7), even with up to $50,000$ traces, the attack fails to differentiate correct weight components from incorrect ones. The absolute correlation values remain relatively stable, indicating that the countermeasure effectively mitigates the CPA attack.
@@ V Discussion
Figure 8: Computation overhead of the proposed protected neural network (MLP with varying numbers of layers) compared to standard shuffling-protected neural network implementations. Each layer contains the same fixed number of neurons, with ReLU as the activation function. The $x-$axis represents the fixed number of neurons per layer.
Figure 9: Computation overhead of the proposed protected neural network (MLP with varying numbers of layers) compared to unprotected implementations. Each layer contains the same fixed number of neurons, with ReLU as the activation function. The $x-$axis represents the fixed number of neurons per layer.
@@ V-A Overheads
Timing. Compared to the original Fisher-Yates shuffling method, the masked shuffling introduces an overhead of 3.38×3.38$\times$3.38 × for $N=1000$ (i.e., with $1000$ input neurons) and 2.48×2.48$\times$2.48 × for $N=100$.
For two-layer network computation with shuffled multiplications, incorporating our proposed shuffling operation (Algorithm 3) introduces an overhead of $4\%$ compared to the standard shuffled computation using the original Fisher-Yates algorithm (Algorithm 1) when the layer has $100$ input neurons and $100$ output neurons, with ReLU as the activation function. When the number of input and output neurons increases to $1000$, the overhead decreases to $0.49\%$. As the number of layers increases, the overhead remains relatively unchanged. Figure 8 presents a plot depicting the overhead as a function of the number of neurons, assuming a fixed number of neurons per layer, for various layer counts in the neural network. The results indicate that the primary determining factor is the number of neurons in each layer. This outcome is expected, as the key variation lies in the shuffling operation preceding the multiplication, addition, and activation functions. Given that modern network architectures often comprise several hundred neurons in each layer, we anticipate the overhead in contemporary neural networks to range between $0.49\%-2\%$.
The decreasing trend in the overhead plot is due to the operations involved in the neuron computation: matrix multiplications and activation function calculation. A naïve matrix multiplication calculation assuming the same number of neurons per layer has a time complexity of $\mathcal{O}(n^{4})$, and an activation function calculation has a time complexity of $\mathcal{O}(n^{2})$ [27]. On the other hand, the Fisher-Yates algorithm has a time complexity of $\mathcal{O}(n)$ [10], and the same holds for Blakley’s modular multiplication and the modulus calculation in Algorithm 3 line 4. Therefore, it is expected that with the increasing number of neurons, the overhead of the proposed protected shuffling diminishes.
Similarly, we evaluated the computational overhead introduced by the proposed protected implementation compared to the unprotected implementation. The results, presented in Figure 9, indicate a relatively stable overhead ranging between $4\%$ and $10\%$.
Memory. The memory overhead compared to Fisher-Yates shuffling-based countermeasures arises from storing the two secret arrays, $S_{1}$ and $S_{2}$. As previously noted, the size of $S_{1}$ and $S_{2}$ is determined by the maximum number of neurons in any layer of the network. Specifically, their size is equal to this maximum number minus two.
@@ VI Conclusion
For future work, it would be interesting to evaluate a combination of different hiding-based countermeasures, such as shuffling and desynchronization, or a combination of hiding and masking. Additionally, it would be of interest to evaluate the shuffling resistance by using different preprocessing and attack analysis methods, such as Scatter [28].
In this appendix, we investigate the impact of varying input values on the performance of the attack and the performace of the attack on other weight values.
A natural question arises as to whether fixing the input values of the remaining six neurons affects the attack results. To address this, we randomly generated six input values and used them as fixed inputs instead of the default value of $0.5$. The attack was then performed on both protected and unprotected implementations.
The results, presented in Figures 10 and 11, correspond to the unprotected and protected implementations, respectively. A comparison with Figures 4 and 5 reveals that both the attack effectiveness and the protection measures exhibit consistent performance across varying input configurations.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 10: CPA attack results for the unprotected implementation, where the inputs to the six neurons (except the first) are fixed to random values. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $1.43$, while the gray lines correspond to incorrect values.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 11: CPA attack results for the protected implementation, where the inputs to the six neurons (except the first) are fixed to random values. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $1.43$, while the gray lines correspond to incorrect values.
To further assess the attack’s effectiveness, we attempted to recover additional weight values. The results, consistent for both protected and unprotected implementations, reinforce our earlier findings. As an example, Figures 12 and 13 illustrate the attack results for recovering the weight associated with the third input neuron and the first output neuron of the first layer. The actual weight value is $0.99$, with the following binary representation: sign bit 00, exponent $126$, first mantissa byte $250$, second mantissa byte $225$, and last seven bits $36$. A comparison with Figures 4 and 5 further validates that the attack remains effective across different scenarios, while the countermeasure continues to provide protection in these cases.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 12: CPA attack results for recovering the weight corresponding to the third input neuron and the first output neuron of the first layer in the unprotected implementation. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $0.99$, while the gray lines correspond to incorrect values.
Refer to caption Refer to caption Refer to caption
(a) sign bit (b) exponent bits (c) first byte of mantissa
Refer to caption Refer to caption
(d) second byte of mantissa (e) last seven bits of mantissa
Figure 13: CPA attack results for recovering the weight corresponding to the third input neuron and the first output neuron of the first layer in the protected implementation. The $y$-axis represents the absolute correlation. The red lines correspond to the correct values associated with the correct weight of $0.99$, while the gray lines correspond to incorrect values.